In the aggregate expenditure model for a closed economy, assuming investment, government spending and taxes are exogenous, if the marginal propensity to consume is 0.8, a simultaneous 50 unit increase in government spending and a 20 unit decrease in investment will change equilibrium income by:


1
Expert's answer
2021-09-27T09:06:45-0400

Let's take Investment"=" I, Government spending"=" G, Consumption"=" C, Tax"=" T and equilibrium income "=" Y

Marginal Propensity to consume(MPS) "=" 0.8

At equilibrium,

"Y=C+I+G"

"Y = C + MPC(Y-T) + I +G."

"Y = C + 0.8(Y) - 0.8(T) + I + G"

"0.2(Y) = C - 0.8(T) + I + G"

Taking differential on both sides, we get

"0.2.\u2206Y = \u2206C - 0.8(\u2206T) + \u2206I + \u2206G"

Since Government Spending increases by 50 and Investment decreases by 20.

Tax and C will remain constant.

Therefore,

"\u2206G = 50, \u2206I = -20, \u2206T = \u2206C\u00b0 = 0"

Then,

"0.2(\u2206Y) = -20 + 50"

"0.2(\u2206Y) = 30"

"∆Y = \frac{30}{0.2}"

"\u2206Y = 150"

Therefore equilibrum income will change by 150 units.